The length of a rectangle is three meters longer than its width. Its area is 40 square meters. What are the rectangle's dimensions?
Answer by richwmiller(17219)

L=w+3
L*W=40
(W+3)*w=40
w^2+3w=40
complete the square
add (3/2)^2 to both sides
w^2+3w+9/4=169/4
(w+3/2)^2=(13/2)^2
w=-3/2+13/2=10/2=5
w=-3/2-13/2=-16/2=-8